# ASTGI: Adaptive Spatio-Temporal Graph Interactions for Irregular Multivariate Time Series Forecasting

### v2

@@ Abstract
Irregular multivariate time series (IMTS) are prevalent in critical domains like healthcare and finance, where accurate forecasting is vital for proactive decision-making. However, the asynchronous sampling and irregular intervals inherent to IMTS pose two core challenges for existing methods: (1) how to accurately represent the raw information of irregular time series without introducing data distortion, and (2) how to effectively capture the complex dynamic dependencies between observation points. To address these challenges, we propose the Adaptive Spatio-Temporal Graph Interaction (ASTGI) framework. Specifically, the framework first employs a Spatio-Temporal Point Representation module to encode each discrete observation as a point within a learnable spatio-temporal embedding space. Second, a Neighborhood-Adaptive Graph Construction module adaptively builds a causal graph for each point in the embedding space via nearest neighbor search. Subsequently, a Spatio-Temporal Dynamic Propagation module iteratively updates information on these adaptive causal graphs by generating messages and computing interaction weights based on the relative spatio-temporal positions between points. Finally, a Query Point-based Prediction module generates the final forecast by aggregating neighborhood information for a new query point and performing regression. Extensive experiments on multiple benchmark datasets demonstrate that ASTGI outperforms various state-of-the-art methods.
22footnotetext: Corresponding author.
@@ 1 Introduction
Irregular Multivariate Time Series Forecasting (IMTSF) is a core problem across numerous critical scientific and engineering domains. Its applications are wide-ranging, from monitoring vital signs in intensive care units to tracking the evolution of environmental indicators in climate science (Yao et al., 2018; Vio et al., 2013; Shukla and Marlin, 2020; Brouwer et al., 2019; Bilos et al., 2021). This type of data is fundamentally defined by two characteristics: 1) intra-series irregularity, where observations of the same variable occur at unequal time intervals, and 2) inter-series asynchrony, where observation timestamps are misaligned across different variables. These traits make it challenging to directly apply existing models designed for regular time series. Furthermore, despite existing research dedicated to addressing irregular time series prediction tasks, current methodological paradigms are commonly constrained by two core challenges, which impede their application potential in complex real-world scenarios.
@@
Figure 1: An illustration of information distortion in mainstream IMTSF paradigms. (a) Raw IMTS. (b) Interpolation-based: Converts irregular series into equally spaced series through numerical interpolation. (c) Time-aligned-based: Maps the observations of all variables to a unified timeline and fills in missing values. (d) Patch-aligned-based: Slices the time series into multiple patches.
@@
The first challenge is how to accurately represent the original information of an irregular time series without introducing information distortion. An accurate representation must preserve the original sampling pattern, as this pattern itself contains key information about the system’s dynamics (Shukla and Marlin, 2020; Li and Marlin, 2020). Conversely, any artificial alteration to the original data structure introduces information distortion, corrupting this intrinsic information and ultimately compromising the model’s predictive performance (Ansari et al., 2023). There are three main representation paradigms, each with its own limitations. 1) Interpolation-based Methods (Figure 1b): This paradigm transforms an irregular series into an equally spaced series through numerical interpolation (Shukla and Marlin, 2021). However, this approach generates artificial data points that were not actually observed, which can introduce bias and distort the original sampling distribution (Shukla and Marlin, 2019; Zhang et al., 2023). 2) Time-aligned-based Methods (Figure 1c): This paradigm maps the observations of all variables to a unified timeline and fills in the missing values (Brouwer et al., 2019). Its main drawback is the loss of precise information about the time intervals between the original observations (Che et al., 2016; Rubanova et al., 2019). 3) Patch-aligned-based Methods (Figure 1d): This paradigm divides the time series into multiple patches (Nie et al., 2023). However, a rigid division may disrupt the continuity of information. Furthermore, intra-patch aggregation can smooth out critical, fine-grained dynamics (Zhang et al., 2024; Luo et al., 2025).
@@
The second challenge lies in how to effectively capture the complex dynamic dependencies among observation points. It is important to note that these two challenges are closely coupled in a progressive manner: accurate representation serves as the prerequisite and foundation for effective dependency modeling. If the raw information is distorted during representation (Challenge 1), the subsequent modeling of dynamic dependencies (Challenge 2) will inevitably be built upon inaccurate data, fundamentally compromising the model’s ability to capture true system dynamics. In many real-world scenarios, the interactions among points in an irregular time series are not static; they evolve dynamically over time. Accurately capturing these dependencies is crucial for understanding the system’s behavior and making precise predictions (Chen et al., 2023; Zhang et al., 2022). However, existing methods generally rely on pre-defined and non-adaptive interaction structures. This means the scope of information exchange between observation points is often limited by a set of fixed, prior rules. For instance, ODE-based Interaction (Figure 2b), information interaction strictly follows the temporal order, meaning an observation can only influence its immediate subsequent state on the timeline and cannot establish a direct connection with more distant historical observations (Rubanova et al., 2019; Brouwer et al., 2019; Schirmer et al., 2022; Bilos et al., 2021). Meanwhile, graph-based methods (Figure 2c) construct connections based on fixed rules, such as belonging to the same time point or the same variable (Yalavarthi et al., 2024; Li et al., 2025). The common limitation of these methods is their inability to dynamically and flexibly identify the truly relevant observation points based on the specific context of each point. As a result, they struggle to capture the deep dynamic correlations that span across time and variables.
@@
To address the above challenges, we propose the Adaptive Spatio-Temporal Graph Interaction (ASTGI) framework (Figure 2d). This framework begins with a Spatio-Temporal Point Representation module that directly encodes each discrete time series observation into a point within a learnable spatio-temporal embedding space. This method operates on the original set of observation points without requiring interpolation or alignment, thus fully preserving the structure and patterns of the raw data and effectively avoiding information distortion, thereby addressing the first challenge. To tackle the second challenge, we then design a Neighborhood-Adaptive Graph Construction module, which adaptively builds a causal graph for each point in the embedding space via a nearest-neighbor search, replacing fixed a priori interaction rules. Subsequently, the Spatio-Temporal Dynamic Propagation module performs iterative information updates on these adaptive graphs, generating messages and calculating interaction weights based on the relative spatio-temporal positions between points. Finally, the framework’s Query Point-based Prediction module yields the prediction by aggregating information from the neighborhood of a new query point and performing regression on it.
@@
  • We conduct extensive experiments on public datasets. The results show that ASTGI outperforms various state-of-the-art baselines.
@@ 2 Related Work
Existing methods for IMTSF can be broadly categorized into two paradigms based on how they handle data irregularity: Structured Representation-based methods and Raw-Data-based methods. The former transforms irregular data into regular structures to utilize standard sequence models, often at the cost of information distortion. The latter models discrete observations directly but typically relies on fixed rules for interaction. Our ASTGI framework falls into the second category but distinguishes itself by employing a fully adaptive graph interaction mechanism to capture dynamic dependencies without information loss.
@@ 2.1 Structured Representation-based Methods
The core idea of these methods is to convert irregular, asynchronous data into a regular format through structural transformation, thereby making it compatible with standard sequence models. This category includes the following mainstream approaches: (1) Interpolation-based Methods: This technique generates new values at missing time points through function fitting to create an equally spaced time series, as adopted in works like mTAN (Shukla and Marlin, 2021). Its main limitation is the introduction of artificial data points that were not actually observed, which can alter the original data distribution and distort its intrinsic dynamic patterns (Shukla and Marlin, 2019; Zhang et al., 2023). (2) Time-Aligned-based Methods: This strategy maps the observations of all variables to a unified global timeline and fills in missing values, but this causes a loss of precise time interval information between original observations, thereby introducing distortion (Che et al., 2016; Rubanova et al., 2019; Brouwer et al., 2019). (3) Patch-Aligned-based Methods: To mitigate the sequence length problem caused by alignment, methods like t-PatchGNN (Zhang et al., 2024) divide the timeline into fixed-size patches for local alignment. However, this aggregation based on a predefined granularity may smooth out or lose critical fine-grained dynamics within each patch (Luo et al., 2025; Nie et al., 2023). In contrast, our ASTGI framework directly represents each discrete observation, fully preserving the original data and fundamentally avoiding the information distortion caused by structural transformations.
@@ 2.2 Raw-Data-based Methods
Unlike the previous category, this paradigm models the set of discrete observation points directly, avoiding the distortion introduced by data structuring. However, it typically relies on predefined, non-adaptive rules to capture the dependencies between points. (1) ODE-based Interaction: Represented by models like Latent-ODE (Rubanova et al., 2019) and NeuralFlows (Bilos et al., 2021), these methods treat the evolution of a time series as a continuous dynamical system. Although they can naturally handle queries at any time point, their inherent Markov assumption strictly confines interactions to temporally adjacent states. This prevents the model from capturing direct long-range dependencies between non-adjacent events (Brouwer et al., 2019; Schirmer et al., 2022). (2) Static Structure Interaction: These methods use Graph Neural Networks (GNNs) to learn the relationships between observation points. However, their graph structure is typically constructed based on fixed heuristic rules (Yalavarthi et al., 2024; Li et al., 2025). Such a static topology is insensitive to the specific data context and cannot adapt as the system state evolves, making it difficult to capture event-driven dynamic associations. Unlike the fixed interaction rules of these methods, ASTGI adaptively constructs an interaction graph for each observation point, enabling it to dynamically capture context-dependent dependencies.
@@ 3 Methodology
${\mathbb{S}}_{\text{hist}}=\{(t_{i},x_{i},c_{i})\in{\mathbb{S}}\mid t_{i}\leq t_{s}\}$
${\mathbb{S}}_{\text{query}}=\{(t_{j},x_{j},c_{j})\in{\mathbb{S}}\mid t_{j}>t_{s}\}$
@@
$\mathcal{F}({\mathbb{S}}_{\text{hist}},{\mathbb{Q}})\rightarrow\hat{{\mathbb{X}}}_{q}$
@@ 3.2 Spatio-Temporal Point Representation
${\bm{p}}_{i}={\bm{e}}_{c_{i}}\oplus{\bm{e}}_{t_{i}}\in\mathbb{R}^{d_{c}+d_{t}}$ (1)
where $\oplus$ denotes vector concatenation. With this, the original set of discrete observations is transformed into a set of spatio-temporal points $\{({\bm{p}}_{i},{\bm{h}}_{i}^{(0)})\mid(t_{i},x_{i},c_{i})\in{\mathbb{S}}_{\text{hist}}\}$. This representation preserves every original observation point in its entirety, thereby avoiding issues common to interpolation or alignment paradigms, such as information distortion and the introduction of artificial data points. The core of this representation lies in the spatio-temporal coordinate space $\mathbb{R}^{d_{c}+d_{t}}$, formed by the channel and time embeddings, which directly serves as a metric space. It is important to emphasize that the spatial dimension here does not refer to physical geographic location. Instead, it is an abstract dimension learned from the data, designed to capture the intrinsic relationships between different variables. The subsequent graph construction is then adaptively defined based entirely on the proximity of points within this learned space.
@@ 3.3 Neighborhood-Adaptive Graph Construction
To overcome the limitations of predefined, static interaction structures, we do not presuppose a global graph structure. Instead, we adaptively construct a directed, and weighted causal graph for each spatio-temporal point $\{({\bm{p}}_{i},{\bm{h}}_{i}^{(0)})\mid(t_{i},x_{i},c_{i})\in{\mathbb{S}}_{\text{hist}}\}$. For clarity, we will refer to an arbitrary spatio-temporal point as point $i$, and one of its neighboring points as point $j$.
@@ 3.3.2 Relation-Aware Scoring
${\bm{r}}_{ij}^{(l)}=({\bm{p}}_{i}-{\bm{p}}_{j})\oplus{\bm{h}}_{i}^{(l)}\oplus{\bm{h}}_{j}^{(l)}\in\mathbb{R}^{(d_{c}+d_{t})+2d_{\text{model}}}$ (2)
@@ 3.4 Spatio-Temporal Dynamic Propagation
On the adaptively constructed graphs, we stack $L$ information propagation layers to update the point features, thereby capturing long-range and complex spatio-temporal dependencies. At layer $l$, the feature ${\bm{h}}_{i}^{(l)}$ of each point $i$ is updated following a message-aggregation-update framework.

### v3

@@ Abstract
Irregular multivariate time series (IMTS) are prevalent in critical domains like healthcare and finance, where accurate forecasting is vital for proactive decision-making. However, the asynchronous sampling and irregular intervals inherent to IMTS pose two core challenges for existing methods: (1) how to accurately represent the raw information of irregular time series without introducing data distortion, and (2) how to effectively capture the complex dynamic dependencies between observation points. To address these challenges, we propose the Adaptive Spatio-Temporal Graph Interaction (ASTGI) framework. Specifically, the framework first employs a Spatio-Temporal Point Representation module to encode each discrete observation as a point within a learnable spatio-temporal embedding space. Second, a Neighborhood-Adaptive Graph Construction module adaptively builds a causal graph for each point in the embedding space via nearest neighbor search. Subsequently, a Spatio-Temporal Dynamic Propagation module iteratively updates information on these adaptive causal graphs by generating messages and computing interaction weights based on the relative spatio-temporal positions between points. Finally, a Query Point-based Prediction module generates the final forecast by aggregating neighborhood information for a new query point and performing forecasting. Extensive experiments on multiple benchmark datasets demonstrate that ASTGI outperforms various state-of-the-art methods.
@@ 1 Introduction
Irregular Multivariate Time Series Forecasting (IMTSF) is a core problem across numerous critical scientific and engineering domains. Its applications are wide-ranging, from monitoring vital signs in intensive care units to tracking the evolution of environmental indicators in climate science (Yao et al., 2018; Vio et al., 2013; Shukla and Marlin, 2020; Brouwer et al., 2019; Bilos et al., 2021; Wang et al., 2023; Shao et al., 2025; Wu et al., 2025c; Cheng et al., 2026; Lu et al., 2011; Yang et al., 2023). This type of data is fundamentally defined by two characteristics: 1) intra-series irregularity, where observations of the same variable occur at unequal time intervals, and 2) inter-series asynchrony, where observation timestamps are misaligned across different variables. These traits make it challenging to directly apply existing models designed for regular time series. Furthermore, despite existing research dedicated to addressing irregular time series prediction tasks (Liu et al., 2026; Qiu et al., 2025a; 2024; Li et al., 2025b; Qiu et al., 2025b), current methodological paradigms are commonly constrained by two core challenges, which impede their application potential in complex real-world scenarios.
@@
Figure 1: An illustration of information distortion in mainstream IMTSF paradigms. (a) Raw IMTS. (b) Interpolation-based: Converts irregular series into equally spaced series through numerical interpolation. Dashed circles denote artificially generated points that were not actually observed. (c) Time-aligned-based: Maps the observations of all variables to a unified timeline. Points on the zero line denote filled placeholders where no real observation exists for that variable. (d) Patch-aligned-based: Slices the time series into multiple patches.
@@
The first challenge is how to accurately represent the original information of an irregular time series without introducing information distortion. An accurate representation must preserve the original sampling pattern, as this pattern itself contains key information about the system’s dynamics (Shukla and Marlin, 2020; Li and Marlin, 2020). Conversely, any artificial alteration to the original data structure introduces information distortion, corrupting this intrinsic information and ultimately compromising the model’s predictive performance (Ansari et al., 2023; Wang et al., 2026b; c; 2025c; 2025b; Qiu et al., 2025c). There are three main representation paradigms, each with its own limitations. 1) Interpolation-based Methods (Figure 1b): This paradigm transforms an irregular series into an equally spaced series through numerical interpolation (Shukla and Marlin, 2021). However, this approach generates artificial data points that were not actually observed, which can introduce bias and distort the original sampling distribution (Shukla and Marlin, 2019; Zhang et al., 2023). 2) Time-aligned-based Methods (Figure 1c): This paradigm maps the observations of all variables to a unified timeline and fills in the missing values (Brouwer et al., 2019; Wang et al., 2025a; Yu et al., 2025b; a). Its main drawback is the loss of precise information about the time intervals between the original observations (Che et al., 2016; Rubanova et al., 2019). 3) Patch-aligned-based Methods (Figure 1d): This paradigm divides the time series into multiple patches (Cirstea et al., 2022; Nie et al., 2023; Wu et al., 2025a). However, a rigid division may disrupt the continuity of information. Furthermore, intra-patch aggregation can smooth out critical, fine-grained dynamics (Zhang et al., 2024b; Luo et al., 2025).
@@
The second challenge lies in how to effectively capture the complex dynamic dependencies among observation points. It is important to note that these two challenges are closely coupled in a progressive manner: accurate representation serves as the prerequisite and foundation for effective dependency modeling. If the raw information is distorted during representation (Challenge 1), the subsequent modeling of dynamic dependencies (Challenge 2) will inevitably be built upon inaccurate data, fundamentally compromising the model’s ability to capture true system dynamics. In many real-world scenarios, the interactions among points in an irregular time series are not static; they evolve dynamically over time. Accurately capturing these dependencies is crucial for understanding the system’s behavior and making precise predictions (Chen et al., 2023b; Zhang et al., 2022; Qiu et al., 2025d; e; Wu et al., 2025b; 2026). However, existing methods generally rely on pre-defined and non-adaptive interaction structures. This means the scope of information exchange between observation points is often limited by a set of fixed, prior rules. For instance, in ODE-based Interaction (Figure 2b), information interaction strictly follows the temporal order, meaning an observation can only influence its immediate subsequent state on the timeline and cannot establish a direct connection with more distant historical observations (Rubanova et al., 2019; Brouwer et al., 2019; Schirmer et al., 2022; Bilos et al., 2021). Meanwhile, graph-based methods (Figure 2c) construct connections based on fixed rules, such as belonging to the same time point or the same variable (Yalavarthi et al., 2024; Li et al., 2025a; Chen et al., 2023a; Shang and Chen, 2024; Shang et al., 2024; 2026). The common limitation of these methods is their inability to dynamically and flexibly identify the truly relevant observation points based on the specific context of each point. As a result, they struggle to capture the deep dynamic correlations that span across time and variables.
@@
To address the above challenges, we propose the Adaptive Spatio-Temporal Graph Interaction (ASTGI) framework (Figure 2d). This framework begins with a Spatio-Temporal Point Representation module that directly encodes each discrete time series observation into a point within a learnable spatio-temporal embedding space. This method operates on the original set of observation points without requiring interpolation or alignment, thus fully preserving the structure and patterns of the raw data and effectively avoiding information distortion, thereby addressing the first challenge. To tackle the second challenge, we then design a Neighborhood-Adaptive Graph Construction module, which adaptively builds a causal graph for each point in the embedding space via a nearest-neighbor search, replacing fixed a priori interaction rules. Subsequently, the Spatio-Temporal Dynamic Propagation module performs iterative information updates on these adaptive graphs, generating messages and calculating interaction weights based on the relative spatio-temporal positions between points. Finally, the framework’s Query Point-based Prediction module yields the prediction by aggregating information from the neighborhood of a new query point and performing forecasting on it.
@@
  • We conduct extensive experiments on public datasets. The results show that ASTGI outperforms various state-of-the-art baselines. All datasets and code are available at https://github.com/decisionintelligence/ASTGI.
@@ 2 Related Work
Deep learning has made impressive progress in natural language processing (NLP), time series analysis, computer vision, and other aspects Zhang et al. (2024a); Yu et al. (2025d; c; e); Zhang et al. (2025a); Chen et al. (2025); Lu et al. (2024a; 2023); Ma et al. (2024a; b); Zhou et al. (2025); Ma et al. (2025d); Wang et al. (2026a). Studies have shown that learned features may perform better than human-designed features Zhang et al. (2025c; b); Qi et al. (2025); Ma et al. (2025c); Lu et al. (2024b; 2025); Ma et al. (2025a; b). Existing deep learning methods for IMTSF can be broadly categorized into two paradigms based on how they handle data irregularity: Structured Representation-based methods and Raw-Data-based methods. The former transforms irregular data into regular structures to utilize standard sequence models, often at the cost of information distortion. The latter models discrete observations directly but typically relies on fixed rules for interaction. Our ASTGI framework falls into the second category but distinguishes itself by employing a fully adaptive graph interaction mechanism to capture dynamic dependencies without information loss.
@@ 2.1 Structured Representation-based Methods
The core idea of these methods is to convert irregular, asynchronous data into a regular format through structural transformation, thereby making it compatible with standard sequence models. This category includes the following mainstream approaches: (1) Interpolation-based Methods: This technique generates new values at missing time points through function fitting to create an equally spaced time series, as adopted in works like mTAN (Shukla and Marlin, 2021). Its main limitation is the introduction of artificial data points that were not actually observed, which can alter the original data distribution and distort its intrinsic dynamic patterns (Shukla and Marlin, 2019; Zhang et al., 2023). (2) Time-Aligned-based Methods: This strategy maps the observations of all variables to a unified global timeline and fills in missing values, but this causes a loss of precise time interval information between original observations, thereby introducing distortion (Che et al., 2016; Rubanova et al., 2019; Brouwer et al., 2019). (3) Patch-Aligned-based Methods: To mitigate the sequence length problem caused by alignment, methods like t-PatchGNN (Zhang et al., 2024b) divide the timeline into fixed-size patches for local alignment. However, this aggregation based on a predefined granularity may smooth out or lose critical fine-grained dynamics within each patch (Luo et al., 2025; Nie et al., 2023). In contrast, our ASTGI framework directly represents each discrete observation, fully preserving the original data and fundamentally avoiding the information distortion caused by structural transformations.
@@ 2.2 Raw-Data-based Methods
Unlike the previous category, this paradigm models the set of discrete observation points directly, avoiding the distortion introduced by data structuring. However, it typically relies on predefined, non-adaptive rules to capture the dependencies between points. (1) ODE-based Interaction: Represented by models like Latent-ODE (Rubanova et al., 2019) and NeuralFlows (Bilos et al., 2021), these methods treat the evolution of a time series as a continuous dynamical system. Although they can naturally handle queries at any time point, their inherent Markov assumption strictly confines interactions to temporally adjacent states. This prevents the model from capturing direct long-range dependencies between non-adjacent events (Brouwer et al., 2019; Schirmer et al., 2022). (2) Static Structure Interaction: These methods use Graph Neural Networks (GNNs) to learn the relationships between observation points. However, their graph structure is typically constructed based on fixed heuristic rules (Yalavarthi et al., 2024; Li et al., 2025a). Such a static topology is insensitive to the specific data context and cannot adapt as the system state evolves, making it difficult to capture event-driven dynamic associations. Unlike the fixed interaction rules of these methods, ASTGI adaptively constructs an interaction graph for each observation point, enabling it to dynamically capture context-dependent dependencies.
@@ 3 Methodology
${\mathbb{S}}_{\text{hist}}=\{(t_{i},x_{i},c_{i})\in{\mathbb{S}}\mid t_{i}\leq t_{s}\},{\mathbb{S}}_{\text{query}}=\{(t_{j},x_{j},c_{j})\in{\mathbb{S}}\mid t_{j}>t_{s}\}.$ (2)
@@
$\mathcal{F}({\mathbb{S}}_{\text{hist}},{\mathbb{Q}})\rightarrow\hat{{\mathbb{X}}}_{q},$ (3)
@@ 3.2 Spatio-Temporal Point Representation
${\bm{p}}_{i}={\bm{e}}_{c_{i}}\oplus{\bm{e}}_{t_{i}}\in\mathbb{R}^{d_{c}+d_{t}},$ (4)
where $\oplus$ denotes vector concatenation. With this, the original set of discrete observations is transformed into a set of spatio-temporal points $\{({\bm{p}}_{i},{\bm{h}}_{i}^{(0)})\mid(t_{i},x_{i},c_{i})\in{\mathbb{S}}_{\text{hist}}\}$. This representation preserves every original observation point in its entirety, thereby avoiding issues common to interpolation or alignment paradigms, such as information distortion and the introduction of artificial data points. The core of this representation lies in the spatio-temporal coordinate space $\mathbb{R}^{d_{c}+d_{t}}$, formed by the channel and time embeddings, which directly serves as a embedding space. It is important to emphasize that the spatial dimension here does not refer to physical geographic location. Instead, it is an abstract dimension learned from the data, designed to capture the intrinsic relationships between different variables. The subsequent graph construction is then adaptively defined based entirely on the proximity of points within this learned space. Notably, the spatio-temporal coordinate ${\bm{p}}_{i}$ remains fixed throughout the subsequent propagation process, serving as a stable positional anchor, while only the feature vector ${\bm{h}}_{i}$ is iteratively refined.
@@ 3.3 Neighborhood-Adaptive Graph Construction
To overcome the limitations of predefined, static interaction structures, we do not presuppose a global graph structure. Instead, we adaptively construct a directed and weighted causal graph for each spatio-temporal point $\{({\bm{p}}_{i},{\bm{h}}_{i}^{(0)})\mid(t_{i},x_{i},c_{i})\in{\mathbb{S}}_{\text{hist}}\}$. For clarity, we will refer to an arbitrary spatio-temporal point as point $i$, and one of its neighboring points as point $j$.
@@ 3.3.2 Relation-Aware Scoring
${\bm{r}}_{ij}^{(l)}=({\bm{p}}_{i}-{\bm{p}}_{j})\oplus{\bm{h}}_{i}^{(l)}\oplus{\bm{h}}_{j}^{(l)}\in\mathbb{R}^{(d_{c}+d_{t})+2d_{\text{model}}}.$ (5)
@@ 3.4 Spatio-Temporal Dynamic Propagation
On the adaptively constructed graphs, we stack $L$ information propagation layers to update the point features, thereby capturing long-range and complex spatio-temporal dependencies. At layer $l$, the feature ${\bm{h}}_{i}^{(l)}$ of each point $i$ is updated following a message-aggregation-update framework. While the graph topology $\mathcal{N}(i)$ remains fixed across layers, the interaction weights $a_{ij}$ are recomputed at each layer based on the updated features, allowing the model to progressively refine its attention over neighbors as representations evolve.